# Efficiently Parameterized Neural Metriplectic Systems

### v2

@@ B.2 Thermoelastic double pendulum
leads to the total energy $E(~{}x)=(1/2m_{1})\mathopen{}\mathclose{{}\left|~{}p_{1}}\right|^{2}+(1/2m_{2})\mathopen{}\mathclose{{}\left|~{}p_{2}}\right|^{2}+E_{1}(~{}x)+E_{2}(~{}x).$
@@
Appendix C Additional experiment: Damped nonlinear oscillator
@@
Here ${}q,~{}p\in\mathbb{R}^{n}$ denote the position and momentum of the oscillator, $S$ is the entropy of a surrounding thermal bath, and the constant parameters $m,\gamma,T$ are the mass, damping rate, and (constant) temperature. This leads to the total energy $E(~{}x)=(1/2m)\mathopen{}\mathclose{{}\left|~{}p}\right|^{2}-k\cos{~{}q}+TS,$ which is readily seen to be constant along solutions ${}x(t)$.
It is now verified that NMS can accurately and stably predict the dynamics of a nonlinear oscillator ${}x=\begin{pmatrix}~{}q&~{}p&S\end{pmatrix}^{\intercal}$ in the case that $n=1,2$, both when the entropy $S$ is observable as well as when it is not. As before, the task considered is prediction in time, although all compared methods NODE, GNODE, and $\text{NMS}_{\text{known}}$ are now trained on full state information from the training interval, and test errors are computed over the full state ${}x$ on the extrapolation interval $(t_{\mathrm{valid}},t_{\mathrm{test}}]$, which is $150\%$ longer than the training interval. In addition, another NMS model, $\text{NMS}_{\mathrm{diff}}$, was trained using only the partial state information ${}x^{o}=\begin{pmatrix}~{}q,~{}p\end{pmatrix}^{\intercal}$and tested under the same conditions, with the initial guess for ${}x^{u}$ generated as in Appendix E. As can be seen in Table 2, NMS is more accurate than GNODE or NODE in both the 1-D and 2-D nonlinear oscillator experiments, improving on previous results by up to two orders of magnitude. Remarkably, NMS produces more accurate entropic dynamics even in the case where the entropic variable $S$ is unobserved during NMS training and observed during the training of other methods. This illustrates another advantage of the NMS approach: because of the reasonable initial data for $S$ produced by the diffusion model, the learned metriplectic system produced by NMS remains performant even when metriplectic governing equations are unknown and only partial state information is observed.
@@
To describe the experimental setup precisely, data is collected from a single trajectory with initial condition as ${}x=(~{}2,~{}0,0)$ following  [13]. The path is calculated at 180,000 steps with a time interval of 0.001, and is then split into training/validation/test sets as before using $t_{\text{train}}=60$, $t_{\text{val}}=90$ and $t_{\text{test}}=180$. Specifications of the networks used for the experiments in Table 2 are:
@@
Appendix D Scaling study
@@
Appendix E Diffusion model for unobserved variables
Recent work in  [16] suggests the benefits of performing time-series generation using a diffusion model. This Appendix describes how this technology is used to generate initial conditions for the unobserved NMS variables in the experiments corresponding to Table 2. More precisely, we describe how to train a conditional diffusion model which generates values for unobserved variables ${}x^{u}$ given values for the observed variables ${}x^{o}$.
@@ Appendix E Diffusion model for unobserved variables
where $t$ is uniformly sampled over $[0,1]$ with an appropriate weight function $\lambda(t):[0,1]\rightarrow\mathbb{R}$. However, using the above formula is computationally prohibitive. Thanks to [29], this loss can be substituted with the following denoising score matching loss:
@@
Since score-based generative models use an affine drift term, the transition kernel $p(\textbf{x}(t)|\textbf{x}(0))$ follows a certain Gaussian distribution [27], and therefore the gradient term ${\nabla}_{\textbf{x}(t)}\log p(\textbf{x}(t)|\textbf{x}(0))$ can be analytically calculated.
@@
As in other diffusion models e.g. [26], a U-net architecture is used, modifying 2-D convolutions to 1-D ones and following the detailed hyperparameters described in [26]. Note the following probability flow ODE seen in [26]:
@@
Although models trained to mimic the probability flow ODE do not match the perofrmance of the forward SDE’s result in the image domain, the authors of [16] observe that the probability flow ODE outperforms the forward SDE in the time-series domain. Therefore, the probability flow ODE is used with the default hyperparameters of [26].

### v3

@@ B.2 Thermoelastic double pendulum
leads to the total energy $E(~{}x)=(1/2m_{1})\mathopen{}\mathclose{{}\left|~{}p_{1}}\right|^{2}+(1/2m_{2})\mathopen{}\mathclose{{}\left|~{}p_{2}}\right|^{2}+E_{1}(~{}x)+E_{2}(~{}x).$ This system is metriplectic with matrix fields ${}L,~{}M$, given by
$~{}L(~{}x)=\begin{pmatrix}~{}0&~{}0&~{}I&~{}0&~{}0\\ ~{}0&~{}0&~{}0&~{}I&~{}0\\ -~{}I&~{}0&~{}0&~{}0&~{}0\\ ~{}0&-~{}I&~{}0&~{}0&~{}0\\ ~{}0&~{}0&~{}0&~{}0&~{}0\end{pmatrix},\qquad~{}M(~{}x)=\begin{pmatrix}~{}0&~{}0&~{}0&~{}0&0&0\\ ~{}0&~{}0&~{}0&~{}0&0&0\\ ~{}0&~{}0&~{}0&~{}0&0&0\\ ~{}0&~{}0&~{}0&~{}0&0&0\\ 0&0&0&0&T_{1}^{-1}T_{2}(~{}x)&-1\\ 0&0&0&0&-1&T_{1}T_{2}^{-1}(~{}x)\end{pmatrix},$
where $T_{i}=\partial_{S_{i}}E_{i}$.
@@
Appendix C Additional figures: Two gas containers and thermoelastic double pendulum
Figure C presents the ground-truth and predicted trajectories of $\dot{q}$ and $\dot{p}$ for the two gas containers example, providing confirmation that the learned metriplectic vector field $\dot{~{}x}$ closely approximates the true metriplectic vector field. The values pictured are obtained by evaluating the different dynamics functions (either the ground truth or those learned by the different parameterization) along the trajectories of $q$ and $p$. Figure C shows that the proposed NMS method produces the most accurate learned velocities $\dot{q},\dot{p}$. The MAE between the ground-truth metriplectic vector field and the predicted ones computes to [0.2652, 0.1355, 0.2285, 0.0546] for [SPNN, GNODE, GFINN, NMS], confirming that NMS yields the lowest error in this quantity. Note that only the observed state variables are included in this computation as no access to the unobserved variables is assumed during training.
(a) $\dot{q}$
(b) $\dot{p}$
Figure 2: The ground-truth and predicted trajectories of $\dot{q}$ and $\dot{p}$ obtained via evaluating the (learned) dynamics functions.
Figure C depicts trajectories of the ground-truth and the predicted position (${}q$), momentum (${}p$), entropy ($S$), and energy ($E$) for the thermoelastic double pendulum example. Figures show that the trajectories produced by NMS follow the ground-truth trajectories closer than those produced by the baseline methods. Figures 3(g) and  3(h) demonstrate that the trajectories of entropy $S$ and energy $E$ produced by GNODE, GFINN, and NMS are thermodynamically consistent (i.e., non-decreasing entropy and conserved energy).
(a) $q_{1x}$
(b) $q_{1y}$
(c) $q_{2x}$
(d) $q_{2y}$
(e) $p_{1x}$
(f) $p_{1y}$
(g) $p_{2x}$
(h) $p_{2y}$
(i) $S$
(j) $E$
Figure 3: The ground-truth and predicted position, momentum, entropy, and energy for the thermoelastic double pendulum. Figures illustrate that the trajectories generated by NMS follow the ground-truth ones closer than those of other methods do.
Appendix D Additional experiment: Damped nonlinear oscillator
@@
Here ${}q,~{}p\in\mathbb{R}^{n}$ denote the position and momentum of the oscillator, $S$ is the entropy of a surrounding thermal bath, and the constant parameters $m,\gamma,T$ are the mass, damping rate, and (constant) temperature. This leads to the total energy $E(~{}x)=(1/2m)\mathopen{}\mathclose{{}\left|~{}p}\right|^{2}-k\cos{~{}q}+TS,$ which is readily seen to be constant along solutions ${}x(t)$. This system has a metriplectic expression in terms of $E,S$ and the matrix fields
$~{}L(~{}x)=\begin{pmatrix}~{}0&-~{}I&0\\ -~{}I&~{}0&0\\ 0&0&0\end{pmatrix},\qquad~{}M(~{}x)=\begin{pmatrix}~{}0&~{}0&0\\ ~{}0&~{}0&0\\ 0&0&\frac{\gamma|~{}q|^{2}}{mT}.\end{pmatrix}$
It is now verified that NMS can accurately and stably predict the dynamics of a nonlinear oscillator ${}x=\begin{pmatrix}~{}q&~{}p&S\end{pmatrix}^{\intercal}$ in the case that $n=1,2$, both when the entropy $S$ is observable as well as when it is not. As before, the task considered is prediction in time, although all compared methods NODE, GNODE, and $\text{NMS}_{\text{known}}$ are now trained on full state information from the training interval, and test errors are computed over the full state ${}x$ on the extrapolation interval $(t_{\mathrm{valid}},t_{\mathrm{test}}]$, which is $150\%$ longer than the training interval. In addition, another NMS model, $\text{NMS}_{\mathrm{diff}}$, was trained using only the partial state information ${}x^{o}=\begin{pmatrix}~{}q,~{}p\end{pmatrix}^{\intercal}$and tested under the same conditions, with the initial guess for ${}x^{u}$ generated as in Appendix H. As can be seen in Table 2, NMS is more accurate than GNODE or NODE in both the 1-D and 2-D nonlinear oscillator experiments, improving on previous results by up to two orders of magnitude. Remarkably, NMS produces more accurate entropic dynamics even in the case where the entropic variable $S$ is unobserved during NMS training and observed during the training of other methods. This illustrates another advantage of the NMS approach: because of the reasonable initial data for $S$ produced by the diffusion model, the learned metriplectic system produced by NMS remains performant even when metriplectic governing equations are unknown and only partial state information is observed.
@@
To describe the experimental setup precisely, data is collected from a single trajectory with initial condition as ${}x=(~{}2,~{}0,0)$ following  Lee et al. (2021). The path is calculated at 180,000 steps with a time interval of 0.001, and is then split into training/validation/test sets as before using $t_{\text{train}}=60$, $t_{\text{val}}=90$ and $t_{\text{test}}=180$. Specifications of the networks used for the experiments in Table 2 are:
@@
Appendix E Additional experiment: Hamiltonian neural networks for two gas container
As briefly mentioned in the main body, from the definition of metriplectic dynamics, restricting $M=0$ leads to Hamiltonian systems. In this section, we present the results of training Hamiltonian systems only on the observable variables $(q,p)$ from the two gas container example problems. We use the same dataset used in Section 5.1 and the same experimental setup described in Section B.1.
For parameterization of the Hamiltonian dynamics, we explicitly consider the canonical Poisson matrix ${}L=~{}J=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ as is standard in Hamiltonian neural networks (HNNs), along with an MLP to parameterize the Hamiltonian function $H=H_{\Theta}(q,p)$, resulting in the dynamics
${\color[rgb]{0,0,0}\dot{~{}x}=\begin{bmatrix}\dot{q}\\ \dot{p}\end{bmatrix}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}\begin{bmatrix}\partial_{q}H\\ \partial_{p}H\end{bmatrix}=~{}L\nabla H.}$ (1)
Figure 4 presents the trajectories of position and momentum predicted by the trained HNN. It essentially shows that imposing an inductive bias (i.e., energy conservation), which does not match with the underlying physical system (i.e., dissipation), results in catastrophic failure in prediction. Figure 4(c) confirms that the learned Hamiltonian is kept constant over the numerical simulation.
(a) $q$
(b) $p$
(c) $\mathcal{H}_{\Theta}$
Figure 4: [Hamiltonian neural networks] The ground-truth and predicted trajectories of position and momentum for the two gas containers example, along with the learned Hamiltonian. Notice that the incorrect inductive bias of conservation leads to catastrophic failure in network prediction.
Appendix F Additional experiment: A damped Thermoelastic rod
As an additional benchmark problem, we consider a damped thermoelastic rod, which discretizes an infinite-dimensional metriplectic system. Consider a 1-dimensional elastic rod, with coordinate $s\in[0,l]$, which evolves following the dynamics defined as:
$\begin{bmatrix}\dot{q}\\ \dot{p}\\ \dot{S}\end{bmatrix}=\begin{bmatrix}\frac{p}{m}\\ V^{\prime}(q)-\gamma\frac{p}{m}\\ \gamma\frac{p^{2}}{m^{2}}\end{bmatrix}=~{}L\nabla E(q,p,e)+~{}M(q,p,e)\nabla S.$ (2)
Here, $V$ denotes a given potential function, $\gamma$ denotes a constant controlling the rate of dissipation, $m=m(s)$ denotes the mass density, and $e=e(s)$ denotes the internal energy representing the conversion of mechanical energy into heat. The energy function is given by
$E(q,p,e)=H(q,p)+S(e)=\int_{0}^{\ell}\mathopen{}\mathclose{{}\left(\frac{p(s)^{2}}{2m(s)}+V(q(s))}\right)+\int_{0}^{\ell}e(s),$ (3)
where $H(q,p)$ denotes the Hamiltonian.
In numerical computation, this continuous system is discretized, leading to $s$ being represented by a vector ${}s\in\mathbb{R}^{N}$, where $N$ is the number of discretization points. This discretization results in a vectorized representation of $q,p$, denoted as ${}q,~{}p\in\mathbb{R}^{N}$, and a discretized version of Eq. (2), where the state variables are $[~{}q,~{}p,S]$. In particular, the resulting metriplectic system is given in terms of $\nabla S=\begin{pmatrix}~{}0&~{}0&1\end{pmatrix}$ and
$~{}L(~{}x)=\begin{pmatrix}~{}0&~{}I&~{}0\\ -~{}I&~{}0&~{}0\\ ~{}0^{\intercal}&~{}0^{\intercal}&0\end{pmatrix},\qquad~{}M(~{}x)=\gamma\begin{pmatrix}~{}0&~{}0&~{}0\\ ~{}0&~{}I&-\frac{~{}p}{m}\\ ~{}0&-\frac{~{}p^{\intercal}}{m}&\mathopen{}\mathclose{{}\left(\frac{\mathopen{}\mathclose{{}\left|~{}p}\right|}{m}}\right)^{2}\end{pmatrix},\qquad\nabla E(~{}x)=\begin{pmatrix}~{}V^{\prime}(~{}q)\\ \frac{~{}p}{m}\\ 1\end{pmatrix}.$
In the experiments, we largely follow the setup used in the two gas container experiments, i.e., testing models’ predictive capability through temporal extrapolation. One simulated trajectory is obtained by solving an IVP with a known thermoelastic rod system, and the trajectory of observed variables is then split into three subsequences, [0, $t_{\text{train}}$], ($t_{\text{train}},t_{\text{val}}]$, and ($t_{\text{val}},t_{\text{test}}]$ for training, validation, and testing with $0<t_{\text{train}}<t_{\text{val}}<t_{\text{test}}$. In the experiment, we set $t_{\text{train}}=30$, $t_{\text{val}}=40$, and $t_{\text{test}}=50$. We train the proposed NMS on spatial discretizations of varying size $N=\{50,75,100\}$, which results in systems of dimension $\{101,151,201\}$, respectively. The performance of NMS is then compared to the performance of GFINNs, which are the current state-of-the-art in metriplectic learning.
Figure F shows the ground truth and predicted trajectories of the first and the last components of $q$ and $p$ for the case of $N=50$, where it can be seen that the trajectories predicted by NMS are significantly better aligned with the ground truth trajectories compared to those of GFINN, particularly in the extrapolation regime ($5040<t<$ 50). Figure F shows the results of training GFINN and NMS on thermoelastic rod systems with varying $N$, where GFINN and NMS instances with comparable model sizes are considered. The results suggest that the proposed NMS is parameter efficient, supporting the theoretical conclusions in Section 3.
(a) $q_{1}$ (first component)
(b) $q_{50}$ (last component)
(c) $p_{1}$ (first component)
(d) $p_{50}$ (last component)
Figure 5: The ground-truth and predicted trajectories of positions and momenta for the thermoelastic rod are depicted with $N=50$.
Figure 6: MSEs (in the $\log_{10}$ scale) of predictions obtained by NMS and GFINN for varying $N$.
Appendix G Scaling study
@@
Appendix H Diffusion model for unobserved variables
Recent work in  Lim et al. (2023) suggests the benefits of performing time-series generation using a diffusion model. This Appendix describes how this technology is used to generate initial conditions for the unobserved NMS variables in the experiments corresponding to Table 2. More precisely, we describe how to train a conditional diffusion model which generates values for unobserved variables ${}x^{u}$ given values for the observed variables ${}x^{o}$.
@@ Appendix H Diffusion model for unobserved variables
where $t$ is uniformly sampled over $[0,1]$ with an appropriate weight function $\lambda(t):[0,1]\rightarrow\mathbb{R}$. However, using the above formula is computationally prohibitive. Thanks to Vincent (2011), this loss can be substituted with the following denoising score matching loss:
@@
Since score-based generative models use an affine drift term, the transition kernel $p(\textbf{x}(t)|\textbf{x}(0))$ follows a certain Gaussian distribution Särkkä and Solin (2019), and therefore the gradient term ${\nabla}_{\textbf{x}(t)}\log p(\textbf{x}(t)|\textbf{x}(0))$ can be analytically calculated.
@@
As in other diffusion models e.g. Song et al. (2020), a U-net architecture is used, modifying 2-D convolutions to 1-D ones and following the detailed hyperparameters described in Song et al. (2020). Note the following probability flow ODE seen in Song et al. (2020):
@@
Although models trained to mimic the probability flow ODE do not match the perofrmance of the forward SDE’s result in the image domain, the authors of Lim et al. (2023) observe that the probability flow ODE outperforms the forward SDE in the time-series domain. Therefore, the probability flow ODE is used with the default hyperparameters of Song et al. (2020).